A sample size of 36 is to be selected from a population that has a mean of µ = 45 and standard deviation s of 10.

  • Find the mean of the sampling distribution
  • Find the standard error of the sampling distribution.
  • What is the probability that this sample mean will be between 40 and 50?
1
Expert's answer
2022-03-25T06:56:45-0400

Mean of sampling distribution = 45


Standard error "=\frac{\sigma}{\sqrt{n}}=\frac{10}{\sqrt{36}}=1.67"


Probability between 40 and 50

"P(40<x<50)=P(\frac{40-45}{\frac{10}{\sqrt{36}}}<Z<\frac{50-45}{\frac{10}{\sqrt{36}}})"

"=P(-3<Z<3)=P(Z<3)-P(Z<-3)"

"=0.9987-0.0013=0.9974"